A man traveling in a train noticed that he passes a telegraphy Pole every 5secs he knows that the pole are 125m apart what speed does he deduce that the train is traveling in kmh
1
Expert's answer
2020-02-14T09:31:27-0500

The speed is


"v=\frac{s}{t}=\frac{125}{5}=25\frac{m}{s}"

"v=(3.6)25=90\frac{km}{h}"